# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ Abstract
Determining the threshold value $\sigma_{*}$ that separates the short-range (SR) and long-range (LR) universality classes in phase transitions remains a controversial issue. While Sak’s criterion, $\sigma_{*}=2-\eta_{\mathrm{SR}}$, has been widely accepted, recent studies of two-dimensional (2D) models with long-range interactions have challenged it. In this work, we focus on the crossover between LR and SR criticality in several classical 2D statistical models, including the XY, Heisenberg, percolation, and Ising models, whose interactions decay as $1/r^{2+\sigma}$. Our previous simulations for the XY, Heisenberg, and percolation models consistently indicate a universal boundary at $\sigma_{*}=2$. Here, we complete the picture by performing large-scale Monte Carlo simulations of the 2D LR-Ising model, reaching lattice sizes up to $L=8192$. By analyzing the Fortuin–Kasteleyn critical polynomial $R_{p}$, the Binder ratio $Q_{m}$, and the anomalous dimension $\eta$, we obtain convergent and self-consistent evidence that the universality class already changes sharply at $\sigma=2$. Taken together, these results establish a unified scenario for LR interacting systems: across all studied models, the crossover from LR to SR universality occurs at $\sigma_{*}=2$.
@@ I Introduction
Long-range (LR) interactions are ubiquitous in physical systems, ranging from dipolar forces in magnetic materials to gravitational interactions in astrophysical systems [campa2014]. These interactions, typically decaying as $1/r^{d+\sigma}$ where $d$ is the spatial dimension and $\sigma$ controls the interaction range, fundamentally alter critical behaviors compared to their short-range (SR) counterparts [spivak2004, lahaye2009, peter2012]. Understanding how LR interactions modify critical phenomena, therefore, remains a central question in statistical and condensed-matter physics. Moreover, recent advances in both theoretical approaches [cpl_42_7_070002, yao2024nonclassical, 2dlrheisen, 2dlrperco] and experimental platforms such as trapped ions and Rydberg atom arrays [lu2012, schauss2012, firstenberg2013, yan2013, aikawa2012, lewis2023] have further stimulated intensive research on this topic.
@@
Despite much progress, several fundamental questions remain unresolved [defenu2023]. Among these, the most debated issue concerns the threshold value of $\sigma$ that separates the LR and SR universality classes. Specifically, for interactions decaying as $1/r^{d+\sigma}$, there exists a critical value $\sigma_{*}$ above which the system’s critical behavior shares the same set of critical exponents as its SR counterpart. In contrast, below $\sigma_{*}$ the system belongs to distinct LR universality classes. The precise value of this threshold has been debated for decades [fisher1972, sak1973, Luijten2002, picco2012, angelini2014, horita2017, cpl_42_7_070002, yao2024nonclassical, 2dlrheisen].
@@
The first systematic investigation of this problem was carried out by Fisher et al. [fisher1972], who employed a second-order $\epsilon$-expansion technique to analyze a field-theoretical description of the LR-O($n$) spin model. This model extends the Ising spin ($s=\pm 1$) to an $n$-component unit vector, with $n=1,2,3$ corresponding to the Ising, XY, and Heisenberg models, respectively. In their work, they proposed three distinct regimes based on the decay exponent $\sigma$: (1) Classical regime ($0<\sigma<d/2$): The critical behavior in this regime is governed by mean-field theory, characterized by Gaussian fixed points; (2) Nonclassical regime ($d/2<\sigma<2$): Gaussian fixed points become unstable, and the critical exponents acquire $\sigma$-dependence. Nonetheless, the anomalous magnetic dimension $\eta$ was conjectured to retain its mean-field value: $\eta=2-\sigma$; (3) SR regime ($\sigma>2$): The LR interactions become irrelevant, and the system exhibits SR critical behavior. According to their analysis, the boundary separating the nonclassical and SR regimes is located at $\sigma_{*}=2$ for all LR-O($n$) models.
@@
However, this scenario implied a discontinuity in the anomalous magnetic dimension $\eta$ at $\sigma=2$, where $\eta$ would abruptly jump from 0 to the SR model’s anomalous magnetic dimension $\eta_{\mathrm{SR}}$, as shown by the sparse dashed line in Fig. 1. While such a discontinuity is not necessarily pathological from the perspective of universality and renormalization group (RG) theory [blumeTheoryFirstOrderMagnetic1966, capelPossibilityFirstorderPhase1966, blumeIsingModelTransition1971], Sak proposed a revised scenario: $\sigma_{*}=2-\eta_{\mathrm{SR}}$ [sak1973], which leads to $\eta=\max(2-\sigma,\eta_{\mathrm{SR}})$. Sak’s prediction thus removes the discontinuity of $\eta$ during the crossover between the LR and SR regimes, as shown by the dense dashed line in Fig. 1. This refinement, now known as Sak’s criterion, has been widely adopted in theoretical studies of LR systems [defenu2023]. Nonetheless, the issue remains debated, with differing viewpoints and no clear consensus in the field-theoretical treatment [YAMAZAKI1977207, YAMAZAKI1978446, vanenter1982, blanchard2013].
@@
From a numerical perspective, the situation is equally challenging and marked by longstanding controversies. In a pioneering work, Luijten and Blöte developed a specialized cluster Monte Carlo (MC) algorithm [luijten1995] for the two-dimensional (2D) LR-Ising model ($\eta_{\rm SR}=1/4$) to investigate the model. By analyzing critical exponents and Binder cumulants [Luijten2002], they found the threshold at $\sigma_{*}=7/4$ supporting Sak’s criterion. However, their data exhibited significant statistical uncertainties near $\sigma=2$ in the 2D case. This issue was later highlighted by Picco [picco2012], who questioned the reliability of Luijten’s conclusions based on the results of the improved numerical techniques and instead proposed a third alternative scenario: As $\sigma\approx 2$, the anomalous dimension $\eta$ smoothly varies from the LR values to the SR value $\eta_{\mathrm{SR}}$, as shown in the blue line of Fig. 1. Blanchard et al. further supported this interpretation with complementary field-theoretical arguments [blanchard2013]. However, the debate persisted when Angelini et al. [angelini2014] challenged Picco’s results from a finite-size scaling (FSS) perspective, employing a double-power scaling approach that led them to question the smooth crossover scenario. Subsequently, Horita et al.[horita2017] introduced the method of self-combined ratios, designed to suppress finite-size corrections more effectively. Their results appeared to support Sak’s criterion once again; however, this technique has a flaw that can potentially mask genuine universal features, as it could artificially compress the difference in the universal values between the nonclassical and SR regimes (see Section III.2 for a detailed discussion).
@@
Figure 1: Sketch of the anomalous dimension $\eta$ as a function of $\sigma$ for four scenarios. For $\sigma>2$, it is uncontroversial that the value belongs to the SR case, as indicated by the black solid line. As for $\sigma\leq 2$, (i) the prediction by Fisher, $\eta=2-\sigma$ [fisher1972], is shown with a sparse dashed line. A jump in $\eta$ occurs at $\sigma=2$, marking the transition point of the universality class at $\sigma_{*}=2$. (ii) Sak’s criterion, $\eta=\max(2-\sigma,\eta_{\rm SR})$ [sak1973], is shown with a dense dashed line. Here the transition point of the universality class is at $\sigma_{*}=2-\eta_{\rm SR}$ (for the Ising model, $\eta_{\rm SR}=1/4$). (iii) Picco proposed a third scenario based on his numerical results [picco2012], where the transition point is at $\sigma_{*}=2$, smoothly connecting to the short-range region, shown by the blue solid line. (iv) We propose a fourth scenario, where although the curve smoothly connects to $\sigma=2$, there is still a jump at $\sigma=2$, indicating that the nature of the phase transition at $\sigma=2$ is already different from the SR case, shown by the red solid line.
@@
Very recently, progress has been made in the study of continuous spin models (O($n$) spin models with $n\geq 2$). The 2D LR dilute XY model [PhysRevB.100.054203] and the standard 2D LR-XY model [cpl_42_7_070002, yao2024nonclassical] consistently indicate that $\sigma_{*}=2$, thus contradicting Sak’s prediction. Meanwhile, these studies did not observe the two phase transitions predicted by Giachetti et al. [giachetti2021, giachetti2022]. Moreover, the 2D LR-Heisenberg model, which exhibits long-range order (LRO) for $\sigma\leq 2$ [2dlrheisen], also supports $\sigma_{*}=2$. Notably, the universal dimensionless ratios and several critical exponents of the 2D LR-percolation model [2dlrperco, grassberger2013], including the anomalous dimension $\eta$, the correlation-length exponent $\nu$ and the shortest-path exponent $d_{\rm min}$, exhibit a clear distinction between $\sigma>2$ and $\sigma\leq 2$, and a discontinuity at $\sigma=2$ has also been observed. These consistent findings across different models strongly suggest $\sigma_{*}=2$, motivating a thorough re-examination of the 2D long-range Ising model, which has historically been the central case of this controversy.
@@
Figure 2: Phase diagrams of the 2D O($n$) spin models: Ising ($n=1$), XY (n=2n=2[cpl_42_7_070002, yao2024nonclassical], and Heisenberg (n=3n=3[2dlrheisen], with long-range (LR) interactions decaying as $1/r^{2+\sigma}$. In all panels, $T_{c}^{\text{CG}}=4/n$ denotes the complete-graph critical temperature [kirkpatrick2017] under the normalization in Eq. (2), and $T_{c}^{\text{NN}}$ indicates the NN limit [PhysRev.65.117, komura2012a]. The classical regime ($\sigma\leq 1$, purple line) follows mean-field behavior. (a) For the Ising case, the system exhibits a second-order transition throughout. The universality class changes at $1<\sigma\leq 2$ (red line), and for $\sigma>2$ (green line), it recovers the SR Ising class. The dashed line indicates that the geometric properties of the low-temperature phase differ between $\sigma<2$ and $\sigma>2$, in the FK representation of the Ising model. (b) For the XY model, the SR regime ($\sigma>2$, brown line) undergoes a Bereinzinskii-Koterlitz-Thouless (BKT) topological phase transition into a quasi-long-range-order (QLRO) phase, while for $1<\sigma\leq 2$, a direct second-order transition into an LRO phase appears. (c) For the Heisenberg model, finite-temperature transitions exist only for $\sigma\leq 2$. All three models consistently exhibit a universality boundary at $\sigma_{*}=2$.
@@
In this work, we provide a unified investigation of the crossover between LR and SR criticality in various 2D statistical models, including the XY, Heisenberg, percolation, and Ising models, with interactions decaying as $1/r^{2+\sigma}$. Building upon our studies on the XY, Heisenberg, and percolation models – each of which consistently suggested a universality boundary at $\sigma_{*}=2$, we now complete the picture by performing large-scale MC simulations of the 2D LR-Ising model. We employ an enhanced variant of the Luijten–Blöte cluster algorithm [luijten1995, luijten1997, michel2019], enabling simulations up to $L=8192$. In particular, we analyze the Ising model from a geometric perspective by measuring the Fortuin–Kasteleyn (FK) critical polynomial $R_{p}$, whose universal value is exactly known for the SR Ising universality class [Jacobsen_2014]. These geometric observables, together with the Binder ratio $Q_{m}$ and the anomalous dimension $\eta$, allow us to determine the universality boundary with high precision and demonstrate that the crossover from LR to SR criticality occurs sharply at $\sigma=2$.
@@
The remainder of this paper is organized as follows. Firstly, in Sec. II, we review our previous studies on the 2D XY [cpl_42_7_070002, yao2024nonclassical], Heisenberg [2dlrheisen], and percolation [2dlrperco] models, which together provide an overview of 2D statistical systems with LR interactions. After that, in Sec. III, we focus on the 2D LR-Ising model and determine its critical points for various $\sigma$ by measuring both the FK critical polynomial $R_{p}$ and the Binder ratio $Q_{m}$. These two observables yield consistent estimates within error bars, as summarized in the phase diagram shown in Fig. 2(a). By performing crossing-point extrapolations, supplemented with least-squares fitting, we obtain the universal values of these observables at criticality, which reveal whether the universality class changes across $\sigma$. Furthermore, the critical exponents are extracted from the FSS of the magnetic susceptibility $\chi$ and the scaled covariance $g^{(x)}_{ER}$, respectively. The resulting quantities, listed in Table 1, consistently demonstrate that the crossover between the LR and SR regimes occurs precisely at $\sigma=2$. In particular, both $R_{p}$ and the anomalous dimension $\eta$ exhibit clear discontinuities at $\sigma=2$, in agreement with the behavior of the LR-percolation model [2dlrperco]. These findings, shown in Figs. 5 and 7, also stand in contrast to earlier numerical studies supporting Sak’s criterion [Luijten2002, horita2017]. Therefore, in Sec. IV, we propose a fourth scenario: although the critical exponents vary smoothly within the LR regime, the crossover at $\sigma_{*}=2$ remains sharp and discontinuous, as illustrated by the red line in Fig. 1. Finally, Sec. V concludes the paper and discusses the broader implications for 2D LR statistical models.
@@ II Overview of the 2D statistical models with long-range interaction
To provide an overview of our findings, Fig. 2 displays the phase diagrams of 2D LR-O($n$) spin models: Ising ($n=1$), XY (n=2n=2[cpl_42_7_070002, yao2024nonclassical], and Heisenberg (n=3n=3[2dlrheisen] models. Furthermore, the phase diagram of the LR-percolation model [2dlrperco] shares the same topology as the LR-Ising case. Also, the phase diagram of the 2D LR uniform forest (UF) model [2dlrheisen], which is a graphical model but can be mapped to a non-Gaussian fermionic theory with non-abelian continuous OSP($1|2$) supersymmetry by generalizing Kirchhoff’s matrix-tree theorem [PhysRevLett.93.080601], has the same topology as that of the Heisenberg model. Remarkably, all these models consistently exhibit a universality boundary at $\sigma_{*}=2$, signaling the crossover between the SR and LR universality classes.
@@
where $\mathbf{S}_{i}$ denotes an $n$-component unit vector at site $i$, and $N=L\times L$ is the total number of sites. The interactions follow the shortest path on its surface, corresponding to the minimum-image convention [frenkel_understanding_2002, christiansen_phase_2019, agrawal_kinetics_2021]. Hence, the summation in Eq. (1) runs over all spin pairs, resulting in N(N1)/2N(N-1)/2 interaction terms. The normalization constant $c(\sigma,L)$ is chosen to satisfy
@@
where the value $4$ corresponds to the number of nearest-neighbor (NN) sites. This choice ensures that the model continuously interpolates between the NN limit ($\sigma\rightarrow\infty$) and the complete-graph (CG) limit ($\sigma\rightarrow-2$) with critical temperature $T_{c}^{\text{CG}}=4/n$ [kirkpatrick2017]. This normalization effectively mitigates finite-size corrections and is asymptotically equivalent to the Ewald summation treatment [frenkel2002, janke2019, agrawal2021]. Importantly, it only rescales the temperature scale and does not alter the physical properties of the model.
@@
For the continuous spin models ($n\geq 2$), the universality can be diagnosed from the intrinsic nature of their phases. As shown in Fig. 2(b) and (c), due to the Mermin–Wagner theorem [mermin_absence_1967], the LRO phase does not exist in the SR case, but appears in the presence of sufficient LR interaction. The rationale is straightforward: the observation of LRO implies that the universality cannot belong to the SR class, unless an extremely nontrivial mechanism intervenes. In the 2D LR-XY [yao2024nonclassical] and Heisenberg [2dlrheisen] models, the systems develop clear LRO through a single phase transition for $\sigma\leq 2$, as evidenced by a finite square magnetization density $\langle M^{2}\rangle>0$ in the limit $L\to\infty$, where $M=L^{-2}\left|\sum_{i}\mathbf{S}_{i}\right|$. To mitigate finite-size corrections and make the existence of LRO more visible, we define the residual squared magnetization as
@@
Moreover, for the continuous spin models, the LRO phase breaks a continuous symmetry, giving rise to Goldstone modes. For $\sigma<2$, the correlation function exhibits an algebraic tail, $g(x)\sim g_{0}+x^{-\eta_{\ell}}$ with $\eta_{\ell}=2-\sigma$ [yao2024nonclassical], leading to the Fourier susceptibility $\chi_{k}=L^{2}\langle M_{k}^{2}\rangle\sim L^{\sigma}$. This scaling of the Fourier susceptibility is the key signature of the LR Goldstone phase, which is clearly observed in [yao2024nonclassical]. For $\sigma=2$, the logarithmic scaling $\chi_{k}\sim L^{2}/\ln(L/L_{0})$ is observed; therefore, we conjecture that the correlation function follows $g(x)\sim g_{0}+1/\ln(x/x_{0})$. Here, $g_{0}$ is the squared magnetization in the thermodynamic limit, $L_{0}$ and $x_{0}$ are non-universal parameters.
Notably, Ref. [PhysRevLett.87.137203] extends the Mermin–Wagner theorem to the LR case, predicting that LRO cannot exist at $\sigma=2$ in 2D for the LR-XY and LR-Heisenberg models. However, our numerical results clearly reveal the presence of LRO and Goldstone modes at $\sigma=2$. Similar behaviors are observed in our Lévy-flight simulations [2dlrheisen], where the nonanalytic propagator $q^{\sigma}$ contributes to the Goldstone mode, and the logarithmic scaling at $\sigma=2$ is also observed. It is also worth emphasizing that the original Mermin–Wagner theorem rules out the spontaneous breaking of continuous symmetry when the second moment of the interaction is finite, i.e., $\sum_{\bm{r}}{\bm{r}}^{2}J(\bm{r})<\infty$ [mermin_absence_1967]. For the interaction $J(r)\propto 1/r^{2+\sigma}$ in 2D, this condition fails at $\sigma=2$, where the summation diverges logarithmically. In any case, the discrepancy between our numerical results and existing theoretical arguments underscores that the nature of the marginal case at $\sigma=2$ deserves further theoretical investigation.

### v2

@@ Abstract
Determining the threshold value $\sigma_{*}$ that separates the short-range (SR) and long-range (LR) universality classes in phase transitions remains a controversial issue. While Sak’s criterion, $\sigma_{*}=2-\eta_{\mathrm{SR}}$, has been widely accepted, recent studies of two-dimensional (2D) models with long-range interactions have challenged it. In this work, we focus on the crossover between LR and SR criticality in several classical 2D statistical models, including the XY, Heisenberg, percolation, and Ising models, whose interactions decay as $1/r^{2+\sigma}$. Our previous simulations for the XY, Heisenberg, and percolation models consistently indicate a universal boundary at $\sigma_{*}=2$. Here, we complete the picture by performing large-scale Monte Carlo simulations of the 2D LR-Ising model, reaching lattice sizes up to $L=8192$. By analyzing the Fortuin-Kasteleyn critical polynomial $R_{p}$, the Binder ratio $Q_{m}$, and the anomalous dimension $\eta$, we obtain convergent and self-consistent evidence that the universality class already changes sharply at $\sigma=2$. Taken together, these results establish a unified scenario for LR interacting systems: across all studied models, the crossover from LR to SR universality occurs at $\sigma_{*}=2$.
@@ I Introduction
Long-range (LR) interactions are ubiquitous in physical systems, ranging from dipolar forces in magnetic materials to gravitational interactions in astrophysical systems [11]. These interactions, typically decaying as $1/r^{d+\sigma}$ where $d$ is the spatial dimension and $\sigma$ controls the interaction range, fundamentally alter critical behaviors compared to their short-range (SR) counterparts [50, 33, 45]. Understanding how LR interactions modify critical phenomena, therefore, remains a central question in statistical and condensed-matter physics. Moreover, recent advances in both theoretical approaches [54, 60, 59, 36] and experimental platforms such as trapped ions and Rydberg atom arrays [37, 49, 17, 58, 3, 34] have further stimulated intensive research on this topic.
@@
Despite much progress, several fundamental questions remain unresolved [16]. Among these, the most debated issue concerns the threshold value of $\sigma$ that separates the LR and SR universality classes. Specifically, for interactions decaying as $1/r^{d+\sigma}$, there exists a critical value $\sigma_{*}$ above which the system’s critical behavior shares the same set of critical exponents as its SR counterpart. In contrast, below $\sigma_{*}$ the system belongs to distinct LR universality classes. The precise value of this threshold has been debated for decades [18, 48, 39, 46, 5, 25, 54, 60, 59, 4].
@@
The first systematic investigation of this problem was carried out by Fisher et al. [18], who employed a second-order $\epsilon$-expansion technique to analyze a field-theoretical description of the LR-O($n$) spin model. This model extends the Ising spin ($s=\pm 1$) to an $n$-component unit vector, with $n=1,2,3$ corresponding to the Ising, XY, and Heisenberg models, respectively. In their work, they proposed three distinct regimes based on the decay exponent $\sigma$: (1) Classical regime ($0<\sigma<d/2$): The critical behavior in this regime is governed by mean-field theory, characterized by Gaussian fixed points; (2) Nonclassical regime ($d/2<\sigma<2$): Gaussian fixed points become unstable, and the critical exponents acquire $\sigma$-dependence. Nonetheless, the anomalous magnetic dimension $\eta$ was conjectured to retain its mean-field value: $\eta=2-\sigma$; (3) SR regime ($\sigma>2$): The LR interactions become irrelevant, and the system exhibits SR critical behavior. According to their analysis, the boundary separating the nonclassical and SR regimes is located at $\sigma_{*}=2$ for all LR-O($n$) models.
@@
However, this scenario implied a discontinuity in the anomalous magnetic dimension $\eta$ at $\sigma=2$, where $\eta$ would abruptly jump from 0 to the SR model’s anomalous magnetic dimension $\eta_{\mathrm{SR}}$, as shown by the sparse dashed line in Fig. 1. While such a discontinuity is not necessarily pathological from the perspective of universality and renormalization group (RG) theory [9, 12, 8], Sak proposed a revised scenario: $\sigma_{*}=2-\eta_{\mathrm{SR}}$ [48], which leads to $\eta=\max(2-\sigma,\eta_{\mathrm{SR}})$. Sak’s prediction thus removes the discontinuity of $\eta$ during the crossover between the LR and SR regimes, as shown by the dense dashed line in Fig. 1. Sak further suggested that the continuous connection between the two regimes, observed at leading order in the $\varepsilon$-expansion, might persist to all perturbative orders. This refinement, now known as Sak’s criterion, has been widely adopted in theoretical studies of LR systems [16]. Nonetheless, the issue remains debated, with differing viewpoints and no clear consensus in the field-theoretical treatment [57, 56, 52, 6].
@@
From a numerical perspective, the situation is equally challenging and marked by longstanding controversies. In a pioneering work, Luijten and Blöte developed a specialized cluster Monte Carlo (MC) algorithm [40] for the two-dimensional (2D) LR-Ising model ($\eta_{\rm SR}=1/4$) to investigate the model. By analyzing critical exponents and Binder cumulants [39], they found the threshold at $\sigma_{*}=7/4$ supporting Sak’s criterion. However, their data exhibited significant statistical uncertainties near $\sigma=2$ in the 2D case. This issue was later highlighted by Picco [46], who questioned the reliability of Luijten’s conclusions based on the results of the improved numerical techniques and instead proposed a third alternative scenario: As $\sigma\approx 2$, the anomalous dimension $\eta$ smoothly varies from the LR values to the SR value $\eta_{\mathrm{SR}}$, as shown in the blue line of Fig. 1. Blanchard et al. further supported this interpretation with complementary field-theoretical arguments [6]. However, the debate persisted when Angelini et al. [5] challenged Picco’s results from a finite-size scaling (FSS) perspective, employing a double-power scaling approach that led them to question the smooth crossover scenario. Subsequently, Horita et al.[25] introduced the method of self-combined ratios, designed to suppress finite-size corrections more effectively. Their results appeared to support Sak’s criterion once again; however, this technique has a flaw that can potentially mask genuine universal features, as it could artificially compress the difference in the universal values between the nonclassical and SR regimes (see Section III.2 for a detailed discussion).
@@
Figure 1: Sketch of the anomalous dimension $\eta$ as a function of $\sigma$ for four scenarios. For $\sigma>2$, it is uncontroversial that the value belongs to the SR case, as indicated by the black solid line. As for $\sigma\leq 2$, (i) the prediction by Fisher, $\eta=2-\sigma$ [18], is shown with a sparse dashed line. A jump in $\eta$ occurs at $\sigma=2$, marking the transition point of the universality class at $\sigma_{*}=2$. (ii) Sak’s criterion, $\eta=\max(2-\sigma,\eta_{\rm SR})$ [48], is shown with a dense dashed line. Here the transition point of the universality class is at $\sigma_{*}=2-\eta_{\rm SR}$ (for the Ising model, $\eta_{\rm SR}=1/4$). (iii) Picco proposed a third scenario based on his numerical results [46], where the transition point is at $\sigma_{*}=2$, smoothly connecting to the short-range region, shown by the blue solid line. (iv) We propose a fourth scenario, where although the curve smoothly connects to $\sigma=2$, there is still a jump at $\sigma=2$, indicating that the nature of the phase transition at $\sigma=2$ is already different from the SR case, shown by the red solid line.
@@
Very recently, progress has been made in the study of continuous spin models (O($n$) spin models with $n\geq 2$). The 2D LR dilute XY model [14] and the standard 2D LR-XY model [54, 60] consistently indicate that $\sigma_{*}=2$, thus contradicting Sak’s prediction. Meanwhile, these studies did not observe the two phase transitions predicted by Giachetti et al. [22, 23]. Moreover, the 2D LR-Heisenberg model, which exhibits long-range order (LRO) for $\sigma\leq 2$ [59], also supports $\sigma_{*}=2$. Notably, the universal dimensionless ratios and several critical exponents of the 2D LR-percolation model [36, 24], including the anomalous dimension $\eta$, the correlation-length exponent $\nu$ and the shortest-path exponent $d_{\rm min}$, exhibit a clear distinction between $\sigma>2$ and $\sigma\leq 2$, and a discontinuity at $\sigma=2$ has also been observed. These consistent findings across different models strongly suggest $\sigma_{*}=2$, motivating a thorough re-examination of the 2D long-range Ising model, which has historically been the central case of this controversy.
@@
Figure 2: Phase diagrams of the 2D O($n$) spin models: Ising ($n=1$), XY (n=2n=2[54, 60], and Heisenberg (n=3n=3[59], with long-range (LR) interactions decaying as $1/r^{2+\sigma}$. In all panels, $T_{c}^{\text{CG}}=4/n$ denotes the complete-graph critical temperature [31] under the normalization in Eq. (2), and $T_{c}^{\text{NN}}$ indicates the NN limit [43, 32]. The classical regime ($\sigma\leq 1$, purple line) follows mean-field behavior. (a) For the Ising case, the system exhibits a second-order transition throughout. The universality class changes at $1<\sigma\leq 2$ (red line), and for $\sigma>2$ (green line), it recovers the SR Ising class. The dashed line indicates that the geometric properties of the low-temperature phase differ between $\sigma<2$ and $\sigma>2$, in the FK representation of the Ising model. (b) For the XY model, the SR regime ($\sigma>2$, brown line) undergoes a Bereinzinskii-Koterlitz-Thouless (BKT) topological phase transition into a quasi-long-range-order (QLRO) phase, while for $1<\sigma\leq 2$, a direct second-order transition into an LRO phase appears. (c) For the Heisenberg model, finite-temperature transitions exist only for $\sigma\leq 2$. All three models consistently exhibit a universality boundary at $\sigma_{*}=2$.
@@
In this work, we provide a unified investigation of the crossover between LR and SR criticality in various 2D statistical models, including the XY, Heisenberg, percolation, and Ising models, with interactions decaying as $1/r^{2+\sigma}$. Building upon our studies on the XY, Heisenberg, and percolation models – each of which consistently suggested a universality boundary at $\sigma_{*}=2$, we now complete the picture by performing large-scale MC simulations of the 2D LR-Ising model. We employ an enhanced variant of the Luijten–Blöte cluster algorithm [40, 38, 42], enabling simulations up to $L=8192$. In particular, we analyze the Ising model from a geometric perspective by measuring the Fortuin–Kasteleyn (FK) critical polynomial $R_{p}$, whose universal value is exactly known for the SR Ising universality class [27]. These geometric observables, together with the Binder ratio $Q_{m}$ and the anomalous dimension $\eta$, allow us to determine the universality boundary with high precision and demonstrate that the crossover from LR to SR criticality occurs sharply at $\sigma=2$.
@@
The remainder of this paper is organized as follows. Firstly, in Sec. II, we review our previous studies on the 2D XY [54, 60], Heisenberg [59], and percolation [36] models, which together provide an overview of 2D statistical systems with LR interactions. After that, in Sec. III, we focus on the 2D LR-Ising model and determine its critical points for various $\sigma$ by measuring both the FK critical polynomial $R_{p}$ and the Binder ratio $Q_{m}$. These two observables yield consistent estimates within error bars, as summarized in the phase diagram shown in Fig. 2(a). By performing crossing-point extrapolations, supplemented with least-squares fitting, we obtain the universal values of these observables at criticality, which reveal whether the universality class changes across $\sigma$. Furthermore, the critical exponents are extracted from the FSS of the magnetic susceptibility $\chi$ and the scaled covariance $g^{(x)}_{ER}$, respectively. The resulting quantities, listed in Table 1, consistently demonstrate that the crossover between the LR and SR regimes occurs precisely at $\sigma=2$. In particular, both $R_{p}$ and the anomalous dimension $\eta$ exhibit clear discontinuities at $\sigma=2$, in agreement with the behavior of the LR-percolation model [36]. These findings, shown in Figs. 5 and 7, also stand in contrast to earlier numerical studies supporting Sak’s criterion [39, 25]. Therefore, in Sec. IV, we propose a fourth scenario: although the critical exponents vary smoothly within the LR regime, the crossover at $\sigma_{*}=2$ remains sharp and discontinuous, as illustrated by the red line in Fig. 1. Finally, Sec. V concludes the paper and discusses the broader implications for 2D LR statistical models.
@@ II Overview of the 2D statistical models with long-range interaction
To provide an overview of our findings, Fig. 2 displays the phase diagrams of 2D LR-O($n$) spin models: Ising ($n=1$), XY (n=2n=2[54, 60], and Heisenberg (n=3n=3[59] models. Furthermore, the phase diagram of the LR-percolation model [36] shares the same topology as the LR-Ising case. Also, the phase diagram of the 2D LR uniform forest (UF) model [59], which is a graphical model but can be mapped to a non-Gaussian fermionic theory with non-abelian continuous OSP($1|2$) supersymmetry by generalizing Kirchhoff’s matrix-tree theorem [13], has the same topology as that of the Heisenberg model. Remarkably, all these models consistently exhibit a universality boundary at $\sigma_{*}=2$, signaling the crossover between the SR and LR universality classes.
@@
where $\mathbf{S}_{i}$ denotes an $n$-component unit vector at site $i$, and $N=L\times L$ is the total number of sites. The interactions follow the shortest path on its surface, corresponding to the minimum-image convention [20, 15, 1]. Hence, the summation in Eq. (1) runs over all spin pairs, resulting in N(N1)/2N(N-1)/2 interaction terms. The normalization constant $c(\sigma,L)$ is chosen to satisfy
@@
where the value $4$ corresponds to the number of nearest-neighbor (NN) sites. This choice ensures that the model continuously interpolates between the NN limit ($\sigma\rightarrow\infty$) and the complete-graph (CG) limit ($\sigma\rightarrow-2$) with critical temperature $T_{c}^{\text{CG}}=4/n$ [31]. This normalization effectively mitigates finite-size corrections and is asymptotically equivalent to the Ewald summation treatment [21, 28, 2]. Importantly, it only rescales the temperature scale and does not alter the physical properties of the model.
@@
For the continuous spin models ($n\geq 2$), the universality can be diagnosed from the intrinsic nature of their phases. As shown in Fig. 2(b) and (c), due to the Mermin–Wagner theorem [41], the LRO phase does not exist in the SR case, but appears in the presence of sufficient LR interaction. The rationale is straightforward: the observation of LRO implies that the universality cannot belong to the SR class, unless an extremely nontrivial mechanism intervenes. In the 2D LR-XY [60] and Heisenberg [59] models, the systems develop clear LRO through a single phase transition for $\sigma\leq 2$, as evidenced by a finite square magnetization density $\langle M^{2}\rangle>0$ in the limit $L\to\infty$, where $M=L^{-2}\left|\sum_{i}\mathbf{S}_{i}\right|$. To mitigate finite-size corrections and make the existence of LRO more visible, we define the residual squared magnetization as
@@
Moreover, for the continuous spin models, the LRO phase breaks a continuous symmetry, giving rise to Goldstone modes. For $\sigma<2$, the correlation function exhibits an algebraic tail, $g(x)\sim g_{0}+x^{-\eta_{\ell}}$ with $\eta_{\ell}=2-\sigma$ [60], leading to the Fourier susceptibility $\chi_{k}=L^{2}\langle M_{k}^{2}\rangle\sim L^{\sigma}$. This scaling of the Fourier susceptibility is the key signature of the LR Goldstone phase, which is clearly observed in [60]. For $\sigma=2$, the logarithmic scaling $\chi_{k}\sim L^{2}/\ln(L/L_{0})$ is observed; therefore, we conjecture that the correlation function follows $g(x)\sim g_{0}+1/\ln(x/x_{0})$. Here, $g_{0}$ is the squared magnetization in the thermodynamic limit, $L_{0}$ and $x_{0}$ are non-universal parameters.
Notably, Ref. [10] extends the Mermin–Wagner theorem to the LR case, predicting that LRO cannot exist at $\sigma=2$ in 2D for the LR-XY and LR-Heisenberg models. However, our numerical results clearly reveal the presence of LRO and Goldstone modes at $\sigma=2$. Similar behaviors are observed in our Lévy-flight simulations [59], where the nonanalytic propagator $q^{\sigma}$ contributes to the Goldstone mode, and the logarithmic scaling at $\sigma=2$ is also observed. It is also worth emphasizing that the original Mermin–Wagner theorem rules out the spontaneous breaking of continuous symmetry when the second moment of the interaction is finite, i.e., $\sum_{\bm{r}}{\bm{r}}^{2}J(\bm{r})<\infty$ [41]. For the interaction $J(r)\propto 1/r^{2+\sigma}$ in 2D, this condition fails at $\sigma=2$, where the summation diverges logarithmically. In any case, the discrepancy between our numerical results and existing theoretical arguments underscores that the nature of the marginal case at $\sigma=2$ deserves further theoretical investigation.